RC meeting and Run Plan, Monday January 9th
From SBS wiki
Return to the RC Meeting Calendar
RC Meeting Information
|Current RC:||Donald Jones|
|Time:||Daily at 3:30PM|
|Location:||Counting House Conference Room|
Remote Connection Details
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
Current Status and Run Plan
- Laser alignment and characterization work continued for part of Saturday and continued early this morning. Radcon support allowed for a paper target to be installed around 7:30AM and Hunter completed his program before 10AM!
- Target cell installation began mid-morning Monday. He3 target, optics ladder and reference cell all installed. Most RTDs installed and connected.
- Jiwan swapped out a faulty NIM crate used in the HCal readout for summing modules. He still has a few dead channels in the trigger that are likely due to faulty pre-amps which he is actively troubleshooting.
- Techs tried flushing the SBS power supply to deal with the overheating issue noticed last week. They met with limited success. Neil Wilson's group will perform a citriflush early Tuesday morning.
- At 5AM Tuesday, Neil Wilson's group will citriflush the SBS power supply for restore cooling circulation. They should be complete by 8AM.
- Tuesday morning target group will install RTD(s) on the target circulation tube, align the cooling jets, pressure test the reference cell to 3 ATM and perform alignment of the target
- Hot checkout Tuesday Day and Lockup Swing
- Wednesday ops is tuning beam into the Hall
Full shifts starting Day Jan 11th. Please sign up
- Jessie reported that things have been progressing well in the hall and that the SBS magnet will be flushed in the morning to prevent overheating. These temperatures are not something we should be monitoring but are part of an automatic shutoff that protects the power supply.
- Bill reported that target installation work was going very well and would be completed before noon Tuesday. Target "Fringe" is now installed.
- Alex reports that the DAQ is ready to go. Sean says that they have run it with GEMs included already.
- Arun reported that the committee investigating the beam strike event that took out the Hall A target has reached the conclusion that the failure to degauss the Moller target magnets was not a significant contributing factor. Daniel Moser is prepared to report to the collaboration/Hall over the next day or two the corrective measures they have taken to assure us that the same thing will not be repeated. The RC will invite him to present at the RC meeting tomorrow afternoon.
- With the possibility of getting beam threaded into the hall as early as Weds, the beam experts will have to be on call.
- Physics starts on Thursday
- Some shifts are still open over the starting few days, so please sign up if you can.
- Ops is delivering fully longitudinal polarization to Hall A until B comes up (earliest possible date for B is Jan 21). We have scheduled a Moller measurement for Monday Jan 16.
Things to follow up on
- Need to follow up with Andrew to see if optics will be updated by Wednesday to allow vertex reconstruction.
- Need a detailed run plan for various optics but in the near term we will be concentrating on just production on He3 since this data point is statistics limited.
- How will we be ramping up to the 45uA we eventually want?